Simplifying: 2(2x+-4)+-3(4x+-1)+2(x+3)=0

Answer by mohd18(3)

so we distribute 2, -3, and 2 to each corresponding values, we get:
4x-8-12x+3+2x+6=0 , we add apples with apples(x) and oranges with oranges(constants), we get: -6x+1 = 0 , -6x=-1, x = (1/6)